OpenStudy (nesha97):

Please explain..... A basketball is dropped from an original height of 9 feet and on each bounce, bounces up 2/3 the distance it fell. How far will it have traveled by the time it hits the ground the fourth time?

OpenStudy (nesha97):

i think the second time is 6...is that right??

OpenStudy (retireed):

Have you tried to draw a picture of the problem? Try to draw it, but I will help if you have a question. I think that once you do, the solution will become obvious. Here is the answer to your first question.... Yes the height after the first bounce is six. I will be waiting for the drawing.

OpenStudy (tarnett1):

(2/3) of 9=6 (2/3) of 6=4 (2/3) 0f 4=2.66 9+6+4+2.66= 21.66' I am pretty sure that this is the answer, but a second opinion wouldn't be a bad idea.

OpenStudy (nesha97):

how would i draw that ?? just with lines??

OpenStudy (retireed):

I do the first bounce if you do the rest. I think @tarnett1 answer is wrong, but it was a good try.

OpenStudy (retireed):